OpenStudy (anonymous):

Which of the following is not a whole number followed by its square? A. 1, 1 B. 6, 36 C. 8, 64 D. 10, 100 E. 11, 144 Can you show me that how to do above?

OpenStudy (anonymous):

E

OpenStudy (anonymous):

1² = 1 * 1 = 1 6² = 6 * 6 = 36 8² = 8 * 8 = 64 10² = 10 * 10 = 100 11² = 11 * 11 = 121